OpenStudy (anonymous):

Which of the following are solutions to the equation below? x2 + 10x + 25 = 12

OpenStudy (anonymous):

For a quadratic equation, \[\Large ax^2 + bx + c = 0\] Use the quadratic formula. \[\Large x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]